OpenStudy (anonymous):

The tangent line at y=f(x) at (-8,-6) passes through the point (1,-3). Compute the following: f(-8) and f'(-8).

OpenStudy (anonymous):

first i would find the equation of my tan line... one way to do this is just use the 2 given points that it passes thru... can you find the line that oes thru (-8,-6) and (1,-3)?????

OpenStudy (anonymous):

The slope of the tangent line is 1/3. But after that what do I do?

OpenStudy (anonymous):

yes the slope is 1/3.... so then we say the line is y=3x+b.... to find b i just plug in a point... to get -6=-8(1/3)+b.... so then what is b?

OpenStudy (anonymous):

y=1/3x+b sorry...

OpenStudy (anonymous):

Is b 9/4?

OpenStudy (anonymous):

not quite... we say that b=(-6+8/3)... so b=-10/3... now we say our tan line is y=1/3x-10/3...

OpenStudy (anonymous):

comp is dying! one sec while i switch out batteries! sorry!

OpenStudy (anonymous):

It's completely fine, thank you so much for your help! Does b=3 then?